# Mixture of In-Context Experts Enhance LLMs’ Long Context Awareness

### v1

@@ Abstract
Many studies have revealed that large language models (LLMs) exhibit uneven awareness of different contextual positions. Their limited context awareness can lead to overlooking critical information and subsequent task failures. While several approaches have been proposed to enhance LLMs’ context awareness, achieving both effectiveness and efficiency remains challenging. In this paper, for LLMs utilizing RoPE as position embeddings, we introduce a novel method called “Mixture of In-Context Experts” (MoICE) to address this challenge. MoICE comprises two key components: a router integrated into each attention head within LLMs and a lightweight router-only training optimization strategy: (1) MoICE views each RoPE angle as an ‘in-context’ expert, demonstrated to be capable of directing the attention of a head to specific contextual positions. Consequently, each attention head flexibly processes tokens using multiple RoPE angles dynamically selected by the router to attend to the needed positions. This approach mitigates the risk of overlooking essential contextual information. (2) The router-only training strategy entails freezing LLM parameters and exclusively updating routers for only a few steps. When applied to open-source LLMs including Llama and Mistral, MoICE surpasses prior methods across multiple tasks on long context understanding and generation, all while maintaining commendable inference efficiency.
@@ 1 Introduction
Although large language models (LLMs) have demonstrated impressive capabilities across diverse NLP tasks, several studies [22, 6] have pointed out that the contextual awareness of LLMs is not as powerful as widely believed, constraining their application in tasks demanding extensive contextual awareness, such as coherent long text generation [38] and Retrieval-Augmented Generation (RAG,  [15, 4, 8]) tasks necessitating in-context retrieval [6]. Liu et al. [22] identified a common issue termed the “lost-in-middle” phenomenon, indicating that LLMs often exhibit a weaker awareness of information situated in the middle of the long context compared to the beginning or end. Chen et al. [6] highlighted challenges arising from a mathematical property of RoPE [29], a wide-used positional embedding in LLMs, which impedes attention to specific positions within the long context. Consequently, if critical information coincides with such positions, task performance suffers.
@@ 2 Background
Positional embedding is crucial for Transformer [33] to perceive sequence order and compensate for the position-agnostic nature of the attention mechanism. In this paper, we mainly focus on LLMs using Rotary Position Embedding (RoPE, [29]) which is the prevalent position embedding in current LLMs. We discuss other position embeddings in Appendix B.
@@
Figure 2: Different RoPE angles $\Theta_{j}$ alter the upper bounds of attention scores between a token and its $x$-distance neighbors. Each angle is distinguished by its own base value $B_{j}$.
LLMs struggle with limited context awareness, significantly impacting their performance in tasks like long-text generation [38], Retrieval-Augmented Generation (RAG, [15, 4, 8]), and multi-turn human-agent interactions [6] involving complex contexts. Liu et al. [22] identified a problem known as “Lost-in-the-Middle,” where LLMs process the beginning and end of the context well but have reduced awareness of the middle. Chen et al. [6] observed that LLMs using RoPE exhibit uneven context awareness, favoring certain positions. Peysakhovich et al. [26] further highlighted that LLMs exhibit variable attention to document-level token segments based on their contextual positions.
@@
According to Chen et al. [6], LLM’s uneven awareness of different contextual positions is due to RoPE’s mathematical characteristics. Within RoPE, the attention score exhibits “waveforms” when retrieving the same token from the context, based on their relative positions. The troughs in these waveforms can impair task performance, especially when critical tokens are situated at these positions during generation. Different RoPE angles produce waveforms with troughs occurring at different positions. These phenomena are depicted in Figure 2. A detailed derivation of the depicted curves in Figure 2 is provided in Appendix A.
@@ 3.1 Architecture
Considering RoPE angles impact how attention heads allocate attention and focus on specific contextual positions, we view each distinct RoPE angle as an in-context expert, in contrast to traditional in-weight experts [14, 17, 10], where the experts are learnable parameter weights. Given that these in-context experts together augment LLMs’ context awareness, we term this method Mixture of In-Context Experts (MoICE, Section 3). Figure 3 illustrates the overview of MoICE. Our proposed MoICE has three major advantages:
@@ 3.2 Router-only training
Given an input sequence, we calculate the negative log-likelihood loss ($\mathcal{L}_{nll}$) for language modeling. During backward propagation, only the parameters of MoICE routers are updated. To mitigate the possibility of a router favoring specific experts disproportionately [14, 34], we incorporate an auxiliary loss $\mathcal{L}_{aux}$ following [14]. Given an input of $T$ tokens and $N$ experts, we calculate the $\mathcal{L}_{lb}$ by the scaled dot-product between frequency vector $\mathbf{F}$ and probability vector $\mathbf{P}$:
@@ 4.2 Long context understanding and generation
We evaluate all these methods alongside our MoICE on two representative open-source LLMs that utilize RoPE for positional embeddings: Llama2-7B-chat [32] with a pre-trained context length of 4,096, and Mistral-7B-Instruct-v0.1 [16]. Mistral-7B employs a sliding window attention (SWA) mechanism with a window size of 4,096 tokens, enabling it to accommodate longer contexts than the default. Therefore, we conduct experiments with a context length of 8,192 on Mistral-7B, using SWA as the exclusive baseline for comparison. For PI and Dynamic NTK, we apply a scaling ratio of 1.5, while for the remaining baselines, we adhere to the hyperparameters specified in their original papers. All methods are tested on a single A800-80G GPU, except for applying AB to Mistral-7B-8k, which needs 2 GPUs due to substantial memory requirements.
@@
We report our experimental results in Table 1. MoICE significantly enhances the overall performance of Llama-2-7B-chat (with p-value < 0.02 in the t-test) in both closed-ended and open-ended tasks. On Mistral, MoICE outperforms all baseline models significantly (p-value < 0.02). Standard fine-tuning degrades the performance of original LLMs, demonstrating catastrophic forgetting and proving that the improvement of MoICE does not stem from more training. These results underscore MoICE’s efficacy in enhancing LLMs’ ability to understand and generate long contexts, both of which require high context awareness. Furthermore, these results underscore the broad applicability of MoICE across different LLMs.
@@ 5 Method analysis
In this section, we delve into a comprehensive analysis of the properties of MoICE. We illustrate how $N$, the total number of in-context experts (Section 5.1), and $K$, the specific number of selected in-context experts (Section 5.2), influence MoICE. We further demonstrate that MoICE is robust to training data (Section 5.3) Additionally, we present a case study demonstrating the dynamic selection of in-context experts for tokens during generation (Section 5.4).
@@ 5.1 The effect of expert total numbers $N$
We investigate the impact of the total number of experts. Employing the search algorithm proposed by Chen et al. [6], we obtain various sets of different sizes, each comprising complementary base values. The searched expert sets are detailed in Appendix C. We apply MoICE to Llama-2-7B-chat and test the model on L-Eval tasks. The results are presented in Table 4. The results of the original Llama are denoted as ($N$=1) in the table. As the table illustrates, MoICE demonstrates improvement to LLMs’ context awareness with increasing $N$, with noticeable improvements even when $N$ is as small as 3. However, as $N$ reaches 9, the average performance is close to $N$=7, indicating a performance plateau. This suggests that having $N$=7 experts is sufficient for general usage.
@@ 6 Conclusion
This work was supported by Boss Zhipin for providing computational resources. We appreciate Ting-En Lin for his discussions on training hyperparameters.
Appendix A Attention waveforms
@@
Appendix B Discussions on more position embeddings
@@
Appendix C Details on expert sets
@@
Appendix D Limitations
In this paper, we introduce a plug-in module called MoICE, which is integrated into the attention heads of open-source LLMs to enhance their context awareness. One limitation is that, due to limited computational resources, we did not investigate the effectiveness of pretraining a language model using the MoICE architecture. Furthermore, our proposed method exploits the potential for context awareness within LLMs, but it does not imbue the models with additional inherent context awareness abilities. Achieving this may necessitate more extensive data to train all model parameters.
Appendix E Broader impacts and safety issues

### v2

@@ Abstract
Many studies have revealed that large language models (LLMs) exhibit uneven awareness of different contextual positions. Their limited context awareness can lead to overlooking critical information and subsequent task failures. While several approaches have been proposed to enhance LLMs’ context awareness, achieving both effectiveness and efficiency remains challenging. In this paper, for LLMs utilizing RoPE as position embeddings, we introduce a novel method called “Mixture of In-Context Experts” (MoICE) to address this challenge. MoICE comprises two key components: a router integrated into each attention head within LLMs and a lightweight router-only training optimization strategy: (1) MoICE views each RoPE angle as an ‘in-context’ expert, demonstrated to be capable of directing the attention of a head to specific contextual positions. Consequently, each attention head flexibly processes tokens using multiple RoPE angles dynamically selected by the router to attend to the needed positions. This approach mitigates the risk of overlooking essential contextual information. (2) The router-only training strategy entails freezing LLM parameters and exclusively updating routers for only a few steps. When applied to open-source LLMs including Llama, Mistral and Qwen, MoICE surpasses prior methods across multiple tasks on long context understanding and generation, all while maintaining commendable inference efficiency. Moreover, we also demonstrate the effectiveness of MoICE in pre-training a language model from scratch.
@@ 1 Introduction
Although large language models (LLMs) have demonstrated impressive capabilities across diverse NLP tasks, several studies [26, 8, 30] have pointed out that the contextual awareness of LLMs is not as powerful as widely believed, constraining their application in tasks demanding extensive contextual awareness, such as in-context learning [27, 45], coherent long text generation [48] and Retrieval-Augmented Generation (RAG,  [19, 6, 10]) tasks necessitating in-context retrieval [8]. Liu et al. [26] identified a common issue termed the “lost-in-middle” phenomenon, indicating that LLMs often exhibit a weaker awareness of information situated in the middle of the long context compared to the beginning or end. Chen et al. [8] highlighted challenges arising from a mathematical property of RoPE [37], a wide-used positional embedding in LLMs, which impedes attention to specific positions within the long context. Consequently, if critical information coincides with such positions, task performance suffers.
@@ 2 Background
Positional embedding is crucial for Transformer [42] to perceive sequence order and compensate for the position-agnostic nature of the attention mechanism. In this paper, we mainly focus on LLMs using Rotary Position Embedding (RoPE, [37]) which is the prevalent position embedding in current LLMs. We discuss other position embeddings in Appendix D.
@@
Figure 2: Different $\Theta_{j}$ alter the upper bounds of attention scores between a token and its $x$-distance neighbors. Each angle is distinguished by its own base value $B_{j}$.
LLMs struggle with limited context awareness, significantly impacting their performance in tasks like long-text generation [48], Retrieval-Augmented Generation (RAG, [19, 6, 10, 39]), and multi-turn human-agent interactions [8] involving complex contexts. Liu et al. [26] identified a problem known as “Lost-in-the-Middle,” where LLMs process the beginning and end of the context well but have reduced awareness of the middle. Chen et al. [8] observed that LLMs using RoPE exhibit uneven context awareness, favoring certain positions. Peysakhovich et al. [33] further highlighted that LLMs exhibit variable attention to document-level token segments based on their contextual positions. Lv et al. [28] observed that language models develop context awareness, especially in their ability to copy, through “grokking [34].” They suggest pre-training models with increased regularization to enhance this capability.
@@
According to Chen et al. [8], LLM’s uneven awareness of different contextual positions is due to RoPE’s mathematical characteristics. Within RoPE, the attention score exhibits “waveforms” when retrieving the same token from the context, based on their relative positions. The troughs in these waveforms can impair task performance, especially when critical tokens are situated at these positions during generation. Different RoPE angles produce waveforms with troughs occurring at different positions. These phenomena are depicted in Figure 2. A detailed derivation of the depicted curves in Figure 2 is provided in Appendix B.
@@ 3.1 Architecture
Considering RoPE angles impact how attention heads allocate attention and focus on specific contextual positions, we view each distinct RoPE angle as an in-context expert, in contrast to traditional in-weight experts [16, 21, 12, 18], where the experts are learnable parameter weights. Given that these in-context experts together augment LLMs’ context awareness, we term this method Mixture of In-Context Experts (MoICE, Section 3). Figure 3 illustrates the overview of MoICE. Our proposed MoICE has three major advantages:
@@ 3.2 Router-only training
Given an input sequence, we calculate the negative log-likelihood loss ($\mathcal{L}_{nll}$) for language modeling. During backward propagation, only the parameters of MoICE routers are updated. To mitigate the possibility of a router favoring specific experts disproportionately [16, 43], we incorporate an auxiliary loss $\mathcal{L}_{aux}$ following [16]. An ablation study on this auxiliary loss is in Table 11. Given an input of $T$ tokens and $N$ experts, we calculate the $\mathcal{L}_{lb}$ by the scaled dot-product between frequency vector $\mathbf{F}$ and probability vector $\mathbf{P}$:
@@ 4.2 Long context understanding and generation
We evaluate all these methods alongside our MoICE on three representative open-source LLMs that utilize RoPE for positional embeddings: Llama2-7B-Chat [41], Mistral-7B-Instruct-v0.1 [20] and Qwen1.5-7B-Chat [2]. Llama2-7B and Qwen1.5-7B support a pre-trained context length of 4,096 and 32,768, respectively. Mistral-7B employs a sliding window attention (SWA) mechanism with a window size of 4,096 tokens, enabling it to accommodate longer contexts than the default. Therefore, we conduct experiments with a context length of 8,192 on Mistral-7B, using SWA as the exclusive baseline for comparison. For PI and Dynamic NTK, we apply a scaling ratio of 1.5, while for the remaining baselines, we adhere to the hyperparameters specified in their original papers. All methods are tested on a single A800-80G GPU, except for applying AB to Mistral-7B-8k, which needs 2 GPUs due to substantial memory requirements.
@@
We report our experimental results in Table 1. MoICE significantly enhances the overall performance of Llama-2-7B-chat (with p-value < 0.02 in the t-test) in both closed-ended and open-ended tasks. On Mistral, MoICE outperforms all baseline models significantly (p-value < 0.02). We also report the mean and standard deviation of MoICE in Table 10. Standard fine-tuning degrades the performance of original LLMs, demonstrating catastrophic forgetting and proving that the improvement of MoICE does not stem from more training. These results underscore MoICE’s efficacy in enhancing LLMs’ ability to understand and generate long contexts, both of which require high context awareness. Furthermore, these results underscore the broad applicability of MoICE across different LLMs.
@@
We also perform further experiments on one additional long context benchmark LongBench [4], which are detailed in Appendix A.
@@ 5 Method analysis
In this section, we delve into a comprehensive analysis of the properties of MoICE. We illustrate how $N$, the total number of in-context experts (Section 5.1), and $K$, the specific number of selected in-context experts (Section 5.2), influence MoICE. We further demonstrate that MoICE is robust to training data (Section 5.3) Additionally, we present a case study demonstrating the dynamic selection of in-context experts for tokens during generation (Section 5.4). Finally, we verify the effectiveness of language model with MoICE architecture in pretraining stage (Section 5.5).
@@ 5.1 The effect of expert total numbers $N$
We investigate the impact of the total number of experts. Employing the search algorithm proposed by Chen et al. [8], we obtain various sets of different sizes, each comprising complementary base values. The searched expert sets are detailed in Appendix E. We apply MoICE to Llama-2-7B-chat and test the model on L-Eval tasks. The results are presented in Table 4. The results of the original Llama are denoted as ($N$=1) in the table. As the table illustrates, MoICE demonstrates improvement to LLMs’ context awareness with increasing $N$, with noticeable improvements even when $N$ is as small as 3. However, as $N$ reaches 9, the average performance is close to $N$=7, indicating a performance plateau. This suggests that having $N$=7 experts is sufficient for general usage.
@@ 5.4 The visualization of dynamic routing states
5.5 Applying MoICE to the pretraining stage
We further evaluate the performance of a language model with MoICE architecture in pretraining stage. Specifically, we train a language model with a Llama architecture of 49M parameters, with and without MoICE respectively. We pretrain a small model from scratch and observe the effectiveness of MoICE. More experimental details can be found in Appendix C. We measure the model’s context awareness on the Key-Value Retrieval [26] task, which uses multiple randomly generated key-value string pairs as prompts to evaluate the model’s ability to extract the correct value corresponding to a given query from the context. One prompt example can be found in Figure 6.
Table 7: Experimental results on the Key-Value Retrieval task [4], We evaluate the model’s awareness of different positions by controlling the position index of the key-value pair corresponding to the query among 10 key-value pairs in the prompt.
Position 1 3 5 7 9
Baseline 0.476 0.324 0.328 0.344 0.502
+ MoICE 0.652 0.762 0.634 0.622 0.814
From the results in Table 7, we can see that our model can significantly increase the contextual capabilities of the pretrained language model, which indicates the potential of scaling up our method in pretraining stage.
@@ 6 Conclusion
This work was supported by the National Natural Science Foundation of China (NSFC Grant No. 62122089), Beijing Outstanding Young Scientist Program NO. BJJWZYJH012019100020098, and Intelligent Social Governance Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China. Ang Lv is supported by the Outstanding Innovative Talents Cultivation Funded Programs 2024 of Renmin University of China.
Appendix A Results on LongBench
LongBench [4] is a benchmark for bilingual, multitask, and comprehensive assessment of long context understanding capabilities of large language models. We choose 16 tasks from LongBench, spanning five long-text application scenarios: Single-Doc QA (NarrativeQA, Qasper, and MultiFieldQA-en), Multi-Doc QA (HotpotQA, 2WikiMQA and Musique), Summarization (GovReport, QMSum and MultiNews), Few-shot Learning (TREC, TriviaQA, SAMSum and LSHT) and Synthetic Tasks (Passage Count, PassageRetrieval-en and PassageRetrieval-zh).
For evaluation, we follow the setup of the LongBench benchmark. The input length of LLama2-7B is set to 4k, Mistral-7B to 8k, and Qwen1.5-7B to 32k. We report the average task scores for each scenario for all methods in Table 8.
Table 8: Experimental results on the LongBench Benchmark [4]. We emphasize the highest score in bold.
Method Single-Doc QA Multi-Doc QA Summarization Few-shot Learning Synthetic Tasks Average
Llama2-7B-Chat [41] 25.54 18.47 23.37 51.78 3.94 24.62
+ PI [7] 23.42 23.73 25.34 51.63 7.63 26.35
+ NTK [15] 24.73 23.67 25.41 51.97 8.33 26.82
+ Ms-PoE [48] 23.68 24.59 25.33 51.66 8.04 26.66
+ AB [8] 27.06 22.94 25.52 52.84 8.62 27.40
+ MoICE (Ours) 26.31 23.70 25.60 52.34 9.71 27.53
Mistral-7B-Instruct-8k [20] 27.20 19.89 24.22 52.41 5.06 25.76
+ PI [7] 30.94 24.94 26.24 49.34 9.35 28.16
+ NTK [15] 30.46 21.21 23.89 52.41 8.44 27.28
+ Ms-PoE [48] 27.90 17.89 20.28 48.59 8.95 24.72
+ AB [8] 29.81 21.95 25.58 54.42 7.89 27.93
+ MoICE (Ours) 31.09 22.98 26.69 55.76 8.02 28.91
Qwen1.5-7B-Chat [2] 34.66 35.91 25.77 56.89 33.83 37.41
+ PI [7] 28.28 17.08 24.60 57.51 32.67 32.03
+ NTK [15] 31.35 23.98 24.95 56.64 32.50 33.88
+ Ms-PoE [48] OOM OOM OOM OOM OOM N/A
+ AB [8] OOM OOM OOM OOM OOM N/A
+ MoICE (Ours) 39.37 37.35 25.81 57.29 34.83 38.93
MoICE consistently shows improved average performance across models with 4k, 8k, and 32k context lengths, surpassing previous competitive approaches.
Appendix B Attention waveforms
@@
Appendix C Experimental details on pretraining
In this section, we provide detailed experimental setup in Section 5.5. The model we has 4 layers, 6 heads per layer, a hidden layer dimension of 512 and an intermediate size of 1280. We train the model using the OpenWebText pretraining [17] dataset. We use four GTX A800-80Gs for training for 600k steps, with a context window of 512. During pretraining, we use a learning rate of 0.004 with a cosine annealing schedule and 6,000 warm-up steps.
Appendix D Discussions on more position embeddings
@@
Appendix E Details on expert sets
@@
Appendix F Limitations
In this paper, we introduce a plug-in module called MoICE, which is integrated into the attention heads of open-source LLMs to enhance their context awareness. One limitation is that, due to limited computational resources, we did not investigate the effectiveness of pretraining a language model with more parameters using the MoICE architecture. Furthermore, our proposed method exploits the potential for context awareness within LLMs, but it does not imbue the models with additional inherent context awareness abilities. Achieving this may necessitate more extensive data to train all model parameters.
Appendix G Broader impacts and safety issues